# Evidence of ferroelectric features in low-density supercooled water from ab initio deep neural-network simulations

### v2

@@ Results
Figure 7: Time series of dipole components from four independent DNN simulations with 1536 water molecules at 235K and 3200 bar. Panels a and c display dipole components for the LDL phase, with spontaneous polarization and collective dipole reorientations visible as a kink in c. Panels b and d show dipole components for the HDL phase, with dipole fluctuations around zero, indicating the absence of spontaneous polarization, and a transition from LDL to HDL in b.
@@
The left two panels display the three dipole components for the LDL trajectories while the right two are those for the HDL. Similar to the situation observed with the small boxes, the large samples also display a tendency to spontaneous polarization - the dipole components take on different magnitudes for the entire 100ns. In Figure 7c, there is also an event similar to the kinks observed in the small box corresponding to collective dipole reorientations. Figure 7b, instead, illustrates a trajectory where the spontaneous polarization irreversibly moves to the HDL phase where all components fluctuate around zero, a feature that is observed for the entire 100ns in Figure 7d. Within the limits of the simulations currently available, we do not observe any striking differences in the magnitude of the polarization fluctuations between LDL and HDL going from the small to large box sizes.
@@
These findings offer new insights and perspectives on behavior of supercooled water close to its critical point. On one side, it is clear that the kinetics within the LDL phase and transitions to the HDL are strongly modulated not only by the mass density order parameter but also by the polarization. This finding should motivate the design of new experimental probes looking specifically at dielectric anomalies under supercooled conditions. It is important to clarify that all our results point to the presence of non-zero polarization within the LDL phase, which we refer to as ferroelectric features. However, this does not imply that the LDL phase is necessarily ferroelectric. An examination of the probability distribution along both the density and polarization (Figure 2) shows that the LDL phase displays some underlying structure which may manifest as a paraelectric glassy phase. Indeed, our data are compatible with a paraelectric phase with large dielectric susceptibility due to large polarization fluctuations. The observed transitions between different polarization states suggest a finite free-energy barrier, but it remains unclear whether this is due to intrinsic paraelectricity or finite-size effects. Additionally, the free-energy landscape might be rough with many minima, or the different observed dipole values may be due to finite-domain effects. Our analysis of different trajectories from larger samples indicates that further work involving enhanced sampling methods is necessary to reach more quantitative conclusions about the nature of the thermodynamics in the LDL polarized phase. Recent theoretical efforts by Pastore et al. [Izzo2024] present an important step to rationalize the thermodynamics associated with the coupling between the density and polarization.
Besides this, our results also offer new challenges from the computational perspective. Specifically, our results are limited to only sub-microsecond timescales and one would need to deploy enhanced sampling techniques along the density and polarization coordinates to adequately make stronger claims about thermodynamic and dynamical quantities. Finally, from a theoretical and perhaps conceptual side, the dynamical features we observed in the LDL phase are akin to what one observes in glassy systems under supercooling [Biroli2011, Berthier2007_1, Berthier2007_2] and it remains an open question to the bearing of our results on the glass-transition of liquid water.
@@
Original data created for the study will be available in a persistent repository upon publication.

### v4

@@ Results
Figure 7: Time series of density and dipole components from four independent DNN simulations with 1536 water molecules at 235K and 3200 bar. Panels a, b, e and f show the density as a function of time for the different simulations. Panels c and g display dipole components for the LDL phase, with spontaneous polarization and collective dipole reorientations visible as a kink in g. Panels d and h show dipole components for the HDL phase, with dipole fluctuations around zero, indicating the absence of spontaneous polarization, and a transition from LDL to HDL in d.
@@
The left panels display the density followed by the three dipole components for the LDL trajectories while the right ones are those for the HDL. Similar to the situation observed with the small boxes, the large samples also display a tendency to spontaneous polarization - the dipole components take on different magnitudes for the entire 100ns. In Figure 7g, there is also an event similar to the kinks observed in the small box corresponding to collective dipole reorientations. Figure 7d, instead, illustrates a trajectory where the spontaneous polarization irreversibly moves to the HDL phase where all components fluctuate around zero, a feature that is observed for the entire 100ns in Figure 7h. Within the limits of the simulations currently available, we do not observe any striking differences in the magnitude of the polarization fluctuations between LDL and HDL going from the small to large box sizes.
@@
These findings offer new insights and perspectives on behavior of supercooled water close to its critical point. On one side, it is clear that the kinetics within the LDL phase and transitions to the HDL are strongly modulated not only by the mass density order parameter but also by the polarization. This finding should motivate the design of new experimental probes looking specifically at dielectric anomalies under supercooled conditions, for example using applied external fields [Cassone2024] which may alter the relative stability of the two phases. It is important to clarify that all our results point to the presence of non-zero polarization within the LDL phase, which we refer to as ferroelectric features. However, this does not imply that the LDL phase is necessarily ferroelectric. An examination of the probability distribution along both the density and polarization (Figure 2) shows that the LDL phase displays some underlying structure which may manifest as a paraelectric glassy phase. Indeed, our data are compatible with a paraelectric phase with large dielectric susceptibility due to large polarization fluctuations. The observed transitions between different polarization states suggest a finite free-energy barrier, but it remains unclear whether this is due to intrinsic paraelectricity or finite-size effects. Additionally, the free-energy landscape might be rough with many minima, or the different observed dipole values may be due to finite-domain effects. Our analysis of different trajectories from larger samples indicates that further work involving enhanced sampling methods is necessary to reach more quantitative conclusions about the nature of the thermodynamics in the LDL polarized phase. There have been several previous theoretical efforts for example, by Men’shikov and Fedichev [Menshikov2011], and Pastore et al. [Izzo2024] to rationalize the thermodynamics associated with the coupling between the density and polarization.
Besides this, our report presents the first quantitative numerical demonstration of this effect from first principles simulations, offering new challenges from the computational perspective. Specifically, our results are limited to only sub-microsecond timescales and one would need to deploy enhanced sampling techniques along the density and polarization coordinates to adequately make stronger claims about thermodynamic and dynamical quantities. Finally, from a theoretical and perhaps conceptual side, the dynamical features we observed in the LDL phase are akin to what one observes in glassy systems under supercooling [Biroli2011, Berthier2007_1, Berthier2007_2] and it remains an open question to the bearing of our results on the glass-transition of liquid water.
@@
The data related to this work, including trained DNN model and the relative dataset, are publicly available for download at https://github.com/cesaremalosso/PNAS-2024-ferroelectric_supercooled_water